The Magellanic Clouds Newsletter

An electronic exchange of Magellanic Clouds information

Edited by: Eva K. Grebel, You-Hua Chu, and Dominik J. Bomans

        mcnews@astro.uiuc.edu
http://www.astro.uiuc.edu/mcnews/MCNews.html
http://www.astro.uni-bonn.de/~mcnews/

No. 14 October 27, 1997

Table of Contents

News and Views  1
Abstracts of 13 refereed papers  2
Abstracts of 2 non-refereed papers 10
Meeting announcement  12

News and Views

Dear Collegues,

To ensure a timely distribution of the increased number of abstracts that we have been receiving, we will try to publish the Magellanic Clouds Newsletter in monthly intervals from now on.

The past issue of MCNews contained the announcement of a multi-orbit HST program to study rich LMC clusters. We encourage you to send in similar communications about other surveys and large-scale projects of special interest to Magellanic Clouds researchers in order to promote information exchange and to stimulate potential collaborations.

The Editors

The Magellanic Clouds Newsletter #14 is available as L^{A}TEXfile, gzipped postscript file, and

HTML document at the URLs http://www.astro.uiuc.edu/mcnews/MCissues.html
and http://www.astro.uni-bonn.de/~mcnews/MCissues.html.

Abstracts of Refereed Papers

Molecular abundances in the Magellanic Clouds: III. A multiline study of five cloud cores

Y.-N. Chin^{1,2}, C. Henkel^{3}, T.J. Millar^{5}, J.B. Whiteoak^{6,7}, and M. Marx-Zimmer^{2}

1 Radioastronomisches Institut der Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany 2 Institute of Astronomy and Astrophysics, Academica Sinica, P.O.Box 1-87 Nankang, 115 Taipei, Taiwan 3 Max-Planck-Institut für Radioastronomie, Auf dem Hügel 69, D-53121 Bonn, Germany 4 European Southern Observatory, Casilla 19001, Santiago 19, Chile 5 Department of Physics, UMIST, P O Box 88, Manchester M60 1QD, United Kingdom 6 Australia Telescope National Facility, Radiophysics Laboratories, P.O. Box 76, Epping, NSW 2121, Australia 7 Paul Wild Observatory, Australia Telescope National Facility, CSIRO, Locked Bag 194, Narrabri NSW 2390, Australia

Detections of CO, CS, SO, C_{2}H, HCO^{+}, HCN, HNC, H_{2}CO, and C_{3}H_{2} are reported from LIRS 36, a star-forming region in the Small Magellanic Cloud. C^{18}O, NO, CH_{3}OH, and most notably CN have not been detected, while the rare isotopes ^{13}CO and, tentatively, C^{34}S are seen. This is so far the most extensive molecular multiline study of an interstellar medium with a heavy element depletion exceeding a factor of four.

The X = N(H_{2})/I_{CO} conversion factor is ≈ 4.8 × 10^{21} cm^{−2} (K km s^{−1})^{−1}, slightly larger than the local Galactic disk value. The CO (1–0) beam averaged column density then becomes N(H_{2}) ≈ 3.7 × 10^{21} cm^{−2} and the density n (H_{2} ≈ 100 cm^{−3}. A comparison with X-values from Rubio et al. (1993a) shows that on small scales (R ≈ 10 pc) X-values are more similar to Galactic disk values than previously anticipated, favoring a neutral interstellar medium of predominantly molecular nature in the cores. The I(^{13}CO)/I(C^{18}O) line intensity ratio indicates an underabundance of ^{12}C^{18}O relative to ^{13}C^{16}O w.r.t. Galactic clouds. I(HCO^{+})/I(HCN) and I(HCN)/I(HNC) line intensity ratios are > 1 and trace a warm (T_{kin} > 10 K) molecular gas exposed to a high ionizing flux. Detections of the CS J=2–1, 3–2, and 5–4 lines imply the presence of a high density core with n (H_{2} = 10^{5} − 10^{7} cm^{−3}. In contrast to star-forming regions in the LMC, the CN 1–0 line is substantially weaker than the corresponding ground rotational transitions of HCN, HNC, and CS. CO, CS, HCO^{+}, and H_{2}CO fractional abundances are a factor ≈ 10 smaller than corresponding values in Galactic disk clouds. Fractional abundances of HCN, HNC, and likely CN are even two orders of magnitude below their ‘normal’, Galactic disk values. The CN/CS abundance ratio is < ∼ 1. Based on chemical model calculations, we suggest that this is because of the small metallicity of the SMC, which affects the destruction of CN but not CS, and because of the high molecular core density which also favors CN destruction.

Accepted by: Astronomy and Astrophysics (Main Journal)

For preprints, contact einmann@biaa21.biaa.sinica.edu.tw
Also available from the URL http://biaa.sinica.edu.tw/einmann/

Observations of ground-state OH in the Large Magellanic Cloud

K. J. Brooks^{1} and J. B. Whiteoak^{2}

1 Astrophysics & Optics, The University of NSW, Sydney 2052, NSW, Australia 2 Australia Telescope National Facility, CSIRO, PO Box 76, Epping, NSW 2121, Australia

We have carried out a series of observations of the 1665- and 1667-MHz transitions of the ^{2}Π_{3/2}, J=3/2 OH ground state towards six selected HII regions in the Large Magellanic Cloud (IRAS 05011-6815 and MRC 0510-689, 0513-694B, 0539-691, 0540-696B, 0540-697A) using the Australia Telescope Compact Array. The study has provided the first accurate positions for known 1665- and 1667-MHz OH masers as well as detecting several new masers. The regions all contain H_{2}O or CH_{3}OH masers but OH masers were detected in only four. The 1.6-GHz continuum emission was also imaged to investigate its spatial relationship to the associated OH maser. Although some masers are close to compact continuum components, in other cases they are near the continuum distribution boundaries and perhaps have been created as a result of the HII region interacting with the surrounding interstellar medium.

Accepted by: Monthly Notices of the Royal Astronomical Society

For preprints, contact kbrooks@wodin.phys.unsw.edu.au
Also available from the URL http://xxx.adelaide.edu.au/find/astro-ph/1/Brooks/0/1/0/past/3/0

The Multi-Phase Medium in the Interstellar Complex N44

S. Kim^{1}, Y.-H. Chu^{2}, L. Staveley-Smith^{3}, and C.R. Smith^{4}

1 Mount Stromlo and Siding Spring Observatories, Weston Creek PO, Canberra, ACT 2611, Australia 2 Astronomy Department, University of Illinois at Urbana-Champaign, Urbana, IL 61801, USA 3 Australia Telescope National Facility, CSIRO, P.O. Box 76, Epping, NSW 2121, Australia 4 Astronomy Department, University of Michigan, Ann Arbor, MI 48109-1090, USA

We have obtained high-resolution H i observations of N44, one of the largest H ii complexes in the Large Magellanic Cloud. The distribution and internal motions of the H i gas show dynamic effects of fast stellar winds and supernova blasts. Numerous H i holes are detected, with the most prominent two corresponding to the optically identified superbubbles Shell 1 and Shell 2. The H i gas associated with Shell 1 shows an expansion pattern similar to that of the ionized gas shell, but the mass and kinetic energy of the H i shell are 3–7 times those of the ionized gas shell. The total kinetic energy of the neutral and ionized gas of Shell 1 is still more than a factor of 5 lower than expected in a pressure-driven superbubble. It is possible that the central OB association was formed in a molecular cloud and a visible superbubble was not fully developed until the ambient molecular gas had been dissociated and cleared away. This hypothesis is supported by the existence of a molecular cloud toward N44 and the fact that the apparent dynamic age of the superbubble Shell 1 is much shorter than the age of its OB association LH47. Accelerated H i gas is detected at the supernova remnant 0523−679. The mass and kinetic energy in the associated H i gas are also much higher than those in the ionized gas of 0523−679. Studies of interstellar gas dynamics using ionized gas alone are clearly inadequate; neutral gas components must be included.

Submitted to: The Astrophysical Journal For preprints, contact sek@mso.anu.edu.au

Also available from the URL http://msowww.anu.edu.au/~sek

The Metallicity and Dust Content of HVC 287.5+22.5+240: Evidence for a Magellanic Clouds Origin

Limin Lu^{1}, Blair D. Savage^{2}, Kenneth R. Sembach^{3}, Bart P. Wakker^{2}, Wallace L.W. Sargent^{1}, and Tom A. Oosterloo^{4}

1 Caltech, 105-24, Pasadena, CA 91125, USA 2 Department of Astronomy, University of Wisconsin, 475 N. Charter Street, Madison, WI 53706, USA 3 Department of Physics and Astronomy, Johns Hopkins University, 3400 N. Charles Street, Baltimore, MD 21218, USA 4 CSIRO, Australia Telescope National Facility, P.O.Box 76, Epping, NSW 2121, Australia

We estimate the abundances of S and Fe in the high velocity cloud HVC 287.5+22.5+240, which has a velocity of +240 km s^{−1} with respect to the local standard of rest and is in the Galactic direction l ∼ 287^{∘} and b ∼ 23^{∘}. The measurements are based on UV absorption lines of these elements in the Hubble Space Telescope spectrum of NGC 3783, a background Seyfert galaxy, as well as new H I 21-cm interferometric data taken with the Australia Telescope. We find S/H=0.25 ± 0.07 and Fe/H=0.033 ± 0.006 solar, with S/Fe=7.6 ± 2.2 times the solar ratio. The S/H value provides an accurate measure of the chemical enrichment level in the HVC, while the super-solar S/Fe ratio clearly indicates the presence of dust, which depletes the gas-phase abundance of Fe. The metallicity and depletion information obtained here, coupled with the velocity and the position of the HVC in the sky, strongly suggest that the HVC originated from the Magellanic Clouds. It is likely (though not necessary) that the same process(es) that generated the Magellanic Stream is responsible for HVC287.5+22.5+240.

Accepted by: The Astronomical Journal

For preprints, contact ll@astro.caltech.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9710045

Hubble Space Telescope Observations of Planetary Nebulae in the

Magellanic Clouds VI: Cycles Four and Five Ultraviolet Spectroscopy Using the Faint Object Spectrograph

E. Vassiliadis^{1}, M.A. Dopita^{2}, R.C. Bohlin^{3}, J.P. Harrington^{4}, H.C. Ford^{5}, S.J. Meatheringham^{2}, P.R. Wood^{2}, T.P. Stecher^{6}, S.P. Maran^{7}

1Instituto de Astrofìsica de Canarias, Via Lactea S/N, E-38200 La Laguna, Tenerife, Spain 2Mount Stromlo and Siding Spring Observatories, Institute of Advanced Studies, The Australian National University, Private Bag, Weston Creek P.O., Weston Creek, ACT 2611, Australia 3Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 4University of Maryland, Department of Astronomy, College Park, MD 20742-2421, USA 5Johns Hopkins University, Department of Physics & Astronomy, Charles & 34th St, Baltimore, MD 21218, USA 6NASA Goddard Space Flight Center, Code 680, Greenbelt, MD 20771, USA 7NASA Goddard Space Flight Center, Code 600, Greenbelt, MD 20771, USA

The Faint Object Spectrograph on board the Hubble Space Telescope is used to obtain spectra in the wavelength range λλ1150–4800 of 3 planetary nebulae in the Small Magellanic Cloud and 10 planetary nebulae in the Large Magellanic Cloud. This sample extends the sample of 12 objects previously observed with HST and reported in Paper III of this series. Observed and dereddened emission line fluxes are presented. Reddening estimates from the He II λ1640/λ4686 flux ratio are generally up to 0.2 dex lower than the reddening derived from the Balmer decrement. Nebular temperatures are estimated from the N^{+} λλ2138, 2142/λλ6548, 6583 flux ratio. Nebular densities are calculated from the O [IV] complex at λ1400, the N [IV] λλ1483, 1487 doublet, and the Si [III] λλ1883, 1892 doublet. Densities calculated using the oxygen lines are comparable to those determined from the optical lines. Densities calculated from the nitrogen lines show a scatter of over 3 dex, which is relatively large compared to the optically derived densities. Three of the five densities derived from the silicon lines are greater than 10000 cm^{−3}. The C^{+2}/O^{+2}, Si^{+2}/C^{+2}, and N^{+2}/O^{+2} ionic abundance ratios are calculated using the available ultraviolet emission features. The C/O and N/O ratios are anticorrelated, which supports the premise that third dredge-up has taken place during the asymptotic giant branch phase. In contrast to Paper III, Type I classification does not imply the presence of Si [III] emission. Three objects show P-Cygni-like line profiles at C IV λλ1548, 1551, indicative of stellar winds.

Accepted by: Astrophysical Journal Supplement For preprints, contact vass@iac.es

The Young Inter-Cloud Population I. Distances and Ages

Serge Demers^{1} and Paolo Battinelli^{2}

1 Université de Montréal, Canada 2 Osservatorio Astronomico di Roma, Italy

Deep BV CCD photometry has been used to characterize the properties of the young stellar population in the outer Small Magellanic Cloud (SMC) Wing and in the bridge between the Magellanic Clouds. Our observations include targets not previously observed and reach two magnitudes fainter than Grondin et al. [AJ, 103, 1234 (1992)] data. They permit to determine, for an assumed abundance, simultaneously the color excess and the distance of each association. Distances decrease from West to East and are compatible with a stellar bridge stretching from the SMC Wing to the Large Magellanic Cloud (LMC). We confirm earlier findings that two associations, close in the sky, lay indeed ∼ 5 kpc away along the line of sight. The ages of the observed main sequences indicate that a burst of star formation occurred between 10 and 25 Myr ago. The star formation started earlier in the SMC Wing where evolved main sequence stars of some 60 Myr are seen. No age gradient is seen across the 10 kpc bridge. Magnitude and color of many more bright main sequence stars would be needed to better define the turn-offs and to determine ages more reliably.

Accepted by: The Astronomical Journal

For paper preprints, contact demers@astro.umontreal.ca

Cepheid Period-Radius and Period-Luminosity Relations and the Distance to the LMC

W.P. Gieren^{1}, P. Fouqué^{2,3}, and M. Gómez^{4}

1 Universidad de Concepción, Departamento de Física, Casilla 4009, Concepción, Chile; wgieren@coma.cfm.udec.cl 2 Observatoire de Paris, Section de Meudon DESPA F-92195 Meudon CEDEX, France 3 European Southern Observatory, Casilla 19001, Santiago 19, Chile; pfouque@eso.org 4 P. Universidad Católica de Chile, Dept. de Astronomía y Astrofísica, Casilla 104, Santiago 22, Chile; mgomez@astro.puc.cl

We have used the infrared Barnes-Evans surface brightness technique to derive the radii and distances of 34 Galactic Cepheid variables. Radius and distance results obtained from both versions of the technique are in excellent agreement. The radii of 28 variables are used to determine the period-radius relation. This relation is found to have a smaller dispersion than in previous studies, and is identical to the period-radius relation found by Laney & Stobie from a completely independent method, a fact which provides persuasive evidence that the Cepheid period-radius relation is now determined at a very high confidence level. We use the accurate infrared distances to determine period-luminosity relations in the V,I,J,H, and K passbands from the Galactic sample of Cepheids. We derive improved slopes of these relations from updated LMC Cepheid samples and adopt these slopes to obtain accurate absolute calibrations of the PL relation. By comparing these relations to the ones defined by the LMC Cepheids, we derive strikingly consistent and precise values for the LMC distance modulus in each of the passbands which yield a mean value of μ_{∘}(LMC) = 18.46 ± 0.02.

By analyzing the observed dispersions of the PL relations defined by the LMC and Galactic samples of Cepheids, we disentangle the contributions due to uncertainties in the reddenings, in distance measurement and due to metallicity effects, and we estimate the intrinsic dispersion of the PL relation with the Wesenheit function. Assuming that the Galactic Cepheid distances are typically accurate to ±3% (as shown in a previous paper), and an intrinsic spread in [Fe/H] of ∼ 0.4 dex among the Cepheids of our sample as obtained by Fry & Carney, the observed dispersion of the Galactic Cepheid PL relation suggests a metallicity dependence of Δμ/Δ[Fe/H] ≈ 0.2, about half the value suggested by Sasselov et al. from EROS data. When we apply this correction, the LMC distance modulus is increased to 18.52±0.06 with most of this uncertainty being due to the adopted metallicity correction.

Our results show that the infrared Barnes-Evans technique is very insensitive to both Cepheid metallicity and adopted reddening, and therefore a very powerful tool to derive accurate distances to nearby galaxies by a direct application of the technique to their Cepheid variables, rather than by comparing PL relations of different galaxies, which introduces much more sensitivity to metallicity and absorption corrections which are usually difficult to determine.

Accepted by: The Astrophysical Journal For preprints, contact P. Fouqué: pfouque@eso.org

Also available from http://xxx.lanl.gov/abs/astro-ph/9710161

Microlensing towards the Small Magellanic Cloud EROS 2 first year survey

N. Palanque-Delabrouille^{1}, C. Afonso^{1}, J.N. Albert^{2}, J. Andersen^{6}, R. Ansari^{2}, É. Aubourg^{1}, P. Bareyre^{1,4}, F. Bauer^{1}, J.P. Beaulieu^{5}, A. Bouquet^{4}, S. Char^{7}, X. Charlot^{1}, F. Couchot^{2}, C. Coutures^{1}, F. Derue^{2}, R. Ferlet^{5}, J.F. Glicenstein^{1}, B. Goldman^{1}, A. Gould^{1,8}, D. Graff^{1}, M. Gros^{1}, J. Haissinski^{2}, J.C. Hamilton^{4}, D. Hardin^{1}, J. de Kat^{1}, É. Lesquoy^{1}, C. Loup^{5}, C. Magneville^{1}, B. Mansoux^{2}, J.B. Marquette^{5}, É. Maurice^{3}, A. Milsztajn^{1}, M. Moniez^{2}, O. Perdereau^{2}, L. Prevot^{3}, C. Renault^{1}, J. Rich^{1}, M. Spiro^{1}, A. Vidal-Madjar^{5}, L. Vigroux^{1}, S. Zylberajch^{1}: The EROS collaboration

1 CEA, DSM, DAPNIA, Centre d’Études de Saclay, F-91191 Gif-sur-Yvette Cedex, France 2 Laboratoire de l’Accélérateur Linéaire, IN2P3 CNRS, Université Paris-Sud, F-91405 Orsay Cedex, France 3 Observatoire de Marseille, 2 place Le Verrier, F-13248 Marseille Cedex 04, France 4 Collège de France, Laboratoire de Physique Corpusculaire, IN2P3 CNRS, 11 place Marcellin Berthelot, F-75231 Paris Cedex, France 5 Institut d’Astrophysique de Paris, INSU CNRS, 98 bis Boulevard Arago, F-75014 Paris, France 6 Astronomical Observatory, Copenhagen University, Juliane Maries Vej 30, DK-2100 Copenhagen, Denmark 7 Universidad de la Serena, Facultad de Ciencias, Departamento de Fisica, Casilla 554, La Serena, Chile 8 Department of Astronomy, Ohio State University, Columbus, OH 43210, USA

We present here an analysis of the light curves of 5.3 million stars in the Small Magellanic Cloud observed by EROS (Expérience de Recherche d’Objets Sombres). One star exhibits a variation that is best interpreted as due to gravitational microlensing by an unseen object. This candidate was also reported by the MACHO collaboration. The Einstein radius crossing time is 123 days (once corrected for blending), corresponding to lensing by a Halo object of 2.6_{−2.3}^{+8.2} M_{⊙}. The maximum magnification is a factor of 2.6. The light curve also displays a periodic modulation with a 2.5% amplitude and a period of 5.1 days. Parallax analysis of the candidate indicates that a halo lens would need to have a mass of at least 1.2 M_{⊙}, although a lens in the SMC could have a mass as low as 0.1 M_{⊙}. We estimate the optical depth for microlensing towards the SMC due to this event to be ∼ 3.3 × 10^{−7}, with an uncertainty dominated by Poisson statistics. We show that this optical depth corresponds to about half that expected for a spherical isothermal Galactic halo comprised solely of such objects, and that it is consistent with SMC self-lensing if the SMC is elongated along the line-of-sight by at least 5 kpc.

Submitted to: Astronomy and Astrophysics

For preprints, contact Nathalie.Delabrouille@cea.fr
Also available from http://xxx.lanl.gov/abs/astro-ph/9710194

LMC Microlenses: Dark or Luminous?

Andrew Gould^{1}

1 Dept. of Astronomy, Ohio State University, Columbus, OH 43210, USA

Zhao has proposed that the microlensing events observed toward the Large Magellanic Cloud (LMC) could be due to faint stars in a dwarf galaxy or tidal debris lying along the line of sight to the LMC. Zaritsky & Lin claim to have detected such a structure which, they believe, could account for most of the observed microlensing optical depth. Here I show that a large-area surface-brightness map made by de Vaucouleurs constrains any such structure to one of four possibilities. Either 1) it does not account for a significant fraction of the observed microlensing, 2) it covers the inner ∼ 3^{∘} of the LMC but does not extend beyond ∼ 5^{∘} from the LMC center, 3) it is smooth on scales of ∼ 15^{∘} in both transverse directions or 4) it has a stellar mass-to-light ratio which exceeds by a factor ∼ 10 that of known stellar populations. The second and third possibilities would not be expected to apply to tidal debris. The last merely rephrases the dark-matter problem in a new form.

Submitted to: The Astrophysical Journal

For preprints, contact gould@astronomy.ohio-state.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9709263

Red Clump Stars as a Tracer of Microlensing Optical Depth

David P. Bennett^{1}

1 University of Notre Dame, Physics Dept., Notre Dame, IN 46556, USA

Zaritsky and Lin have recently suggested that the color magnitude diagram of the Large Magellanic Cloud (LMC) contains evidence of foreground red clump stars. They interpret this as evidence of tidal debris or a dwarf galaxy at a distance of ∼ 35 kpc which may be responsible for the large gravitational microlensing optical depth observed by the MACHO Collaboration. I derive a relationship between the microlensing optical depth of such a foreground population and the observed density of foreground red clump stars. Recent observational determinations for Pop I and Pop II stellar mass functions are used to show that the surface density of foreground red clump stars claimed by Zaritsky and Lin implies a microlensing optical depth in the range τ_{fg} = 0.8 − 3.6 × 10^{−8} which is only 3-13% of τ_{LMC} as determined by the MACHO Collaboration. If the foreground population has a similar star formation history to the LMC, then the implied τ_{fg} is only 3-4% of τ_{LMC}.

Submitted to: Astrophysical Journal Letters For preprints, contact bennett@nd.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9710133

A Test For An Intervening Stellar Population As The Origin of Microlensing Events Toward the Large Magellanic Cloud

Eyal Maoz^{1}

1 Astronomy Department, University of California, Berkeley, CA 94720, USA

Zaritsky and Lin (1997; ZL97) have recently presented observations which can be interpreted as evidence for an intervening stellar population along the line-of-sight to the LMC, at a distance of about 34 kpc. The evidence is based on detected populations of red-clump and lower main-sequence stars which appear brighter than the principal distributions by about 0.9 magnitude, and which have similar fractional surface densities of ∼6%. ZL97 suggest that the intervening population may belong to a low surface-brightness dwarf galaxy, or to the tidal debris tail of a disrupted galaxy. They derive an optical depth for gravitational lensing through the intervening system which is consistent with the observed value, thus suggesting that MACHOs comprise at most a small fraction of the Galactic dark halo.

Such interpretation of the microlensing events toward the LMC has major implications for the nature of the dark matter, and thus should be tested in every conceivable way. We suggest a test which is based on the microlensing data themselves: if the observed events arise in the ZL97 intervening population, then all the lenses are at nearly the same distance, and have nearly the same transverse velocity. Thus, the predicted distribution of event durations should directly reflect the stellar mass function in the intervening population, which is probably not significantly different from that in the Galactic disk.

We find that the first eight MACHO events toward the LMC are as consistent with a standard halo distribution of 0.5 M_{⊙} MACHOs (best fit), as they are with an intervening population having a transverse speed of ∼ 240 km s^{−1} and a Miller-Scalo stellar IMF. Monte Carlo simulations reveal that in order to completely rule out either the standard halo distribution of MACHOs or the ZL97 intervening population, about 40 events will be required if the IMF is rising below the H-burning mass limit, and ≈140 if it is falling. If the intervening population hypothesis is correct then: (a) the currently observed range of event durations directly implies a lens mass range of ≈ 0.06–0.8M_{⊙}, which brackets the mass of ≈0.2 M_{⊙} where many investigations claim that the stellar mass function peaks, and (b) the event distribution provides valuable information on the low-mass end of the stellar IMF.

Submitted to: The Astrophysical Journal (Letters)

For preprints, contact maoz@astro.berkeley.edu
Also available from http://xxx.lanl.gov/abs/astro-ph/9710099

Red Clump Morphology as Evidence Against an Intervening Stellar Population as the Primary Source of Microlensing Toward the LMC

J.P. Beaulieu and Penny D. Sackett

Kapteyn Astronomical Institute 9700 AV Groningen, The Netherlands

We examine the morphology of the color-magnitude diagram (CMD) for core helium-burning (red clump) stars to test the recent suggestion by Zaritsky & Lin (1997) that an extension of the red clump in the Large Magellanic Cloud (LMC) toward brighter magnitudes is due to an intervening population of stars that is responsible for a significant fraction of the observed microlensing toward the LMC. Using our own high resolution CCD photometry of several fields across the LMC, we confirm the presence of this additional red clump feature, but conclude that it is caused by stellar evolution rather than a foreground population. We do this by demonstrating that the feature (1) is present in all our LMC fields and in some fields extends to slightly fainter as well as brighter magnitudes compared to the primary clump, (2) is in precise agreement with the location of the blue loops in the isochrones of intermediate age red clump stars with the metallicity and age of the LMC, and (3) is present in the Hipparcos CMD for the solar neighborhood where an intervening population cannot be invoked. Finally, we show that current understanding of stellar evolution and LMC star formation history is consistent with the features seen not only in our data, but also the LMC CMD of Zaritsky & Lin.

Submitted to: The Astronomical Journal

For preprints, contact beaulieu@astro.rug.nl,psackett@astro.rug.nl
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9710156

Probes for Nearby Galaxies

David Crampton^{1}, G. Gussie^{1}, A.P. Cowley^{2} and P.C. Schmidtke^{2}

1 Dominion Astrophysical Observatory, National Research Council of Canada, Victoria, B.C. V8X 4M6, Canada 2 Department of Physics & Astronomy, Arizona State University, Tempe, AZ, 85287-1504, USA

Data are given for 16 extragalactic objects (predominantly AGN) behind the Magellanic Clouds and for 146 quasar candidates behind the nearby galaxies NGC 45, 185, 253, 2366, 2403 and 6822, IC 1613, M31 and M33. The Magellanic Cloud objects were identified by their X-ray emission, and precise optical and X-ray positions and optical photometry and spectra are presented for all of these. The quasar candidates surrounding the other nearby galaxies were identified through a CFHT slitless spectral survey. Although redshifts for only eight of these candidates have been obtained, previous observations indicate that the majority are likely to be quasars. A subsample of 49 of the brighter objects could confidently be used, in addition to the Magellanic Cloud sources, as probes of the gas in nearby galaxies for rotation curve studies, for studies of their halos, for comparison with higher redshift QSO absorption lines, or as references for proper motion studies.

Accepted by: The Astronomical Journal For preprints, contact crampton@dao.nrc.ca

Also available from the URL http://www.hia.nrc.ca/DAO/SCIENCE/science.html

Abstracts of Non-Refereed Papers

A New Stellar Generation in 30 Doradus

Nolan R. Walborn and Rodolfo H. Barbá

Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA

The increasing evidence for current star formation within the 30 Doradus Nebula is reviewed and illustrated. A new stellar generation appears to have been triggered by the massive central cluster (R136), in and just beyond the surrounding, curved nebular filaments. These filaments appear to be ‘star-formation fronts’ moving into the dense dust clouds beyond. The evidence consists of IR continuum and H_{2} sources, an H_{2}O maser, early O dwarfs embedded in dense nebular knots, and intricate bright and dark nebular structures in HST/WFPC2 images which are signatures of starformation activity. 30 Doradus is the nearest and hence most highly resolved extragalactic starburst; an understanding of its complex structure and evolution is essential for interpreting the integrated properties of more distant, unresolved starburst regions. Two-stage starbursts, in which an initial massive, centrally condensed generation subsequently triggers a secondary event around its periphery, may be characteristic of star formation on this scale.

To appear in Boulder-Munich II: Properties of Hot, Luminous Stars (ASP Conf. Series, Vol. XXX), ed. I. D. Howarth (ASP, San Francisco), 1998 For preprints, contact walborn@stsci.edu

The Initial Mass Function of Massive Stars in the Local Group

Philip Massey^{1}

1 KPNO/NOAO, P.O. 26732, Tucson, AZ, 85726-6732, USA

The present-day mass function of massive stars can be measured directly by counting stars as a function of bolometric luminosity; if we understand the star forming history of a region, we can convert this to an initial mass function (IMF). We discuss the complications inherent in this process: because of the high effective temperatures characteristic of massive stars, the bolometric corrections are a very steep function of T_{eff}, but at the same time the near-UV and optical colors are quite insensitive to effective temperatures for T_{eff} >30,000^{∘}. Hence such diagnostic tools as the visual luminosity function are nearly useless as a measure of the IMF. However, the combination of spectroscopy with photometry provides the means of measuring the IMF of massive stars in the differing physical environments of Local Group galaxies. We find that the IMF slope is essentially Salpeter (Γ ∼ −1.3) in the clusters and associations of the Magellanic Clouds and Milky Way (a factor of 10 spread in metallicity), and in both sparse OB associations and the “super star cluster” R136 (a factor of 1000 in the surface densities of stars). Nevertheless, the significant population of high mass stars that appear to be born in the field—presumably of very modest star-forming events—show a considerably steeper slope, with Γ ∼ −4. The evidence suggests that there is no “upper mass cutoff” to the IMF, or at least none that have been detected observationally: the richer and younger a star-forming region, the higher the highest mass star we see, and it is statistics, rather than physics, that limits the mass we see. We discuss the star formation history of OB associations, and find that the formation of intermediate mass stars commences before the formation of high mass stars by several million years. The vast majority of the high mass stars form over a very short period (Δτ <1–2 Myr). Once the high mass stars form, the production of intermediate mass stars usually continues, until the gas is cleared by subsequent supernovae. There are exceptions, however: in the case of very rich clusters (such as R136), the actions of stellar winds is apparently sufficient to clear out the gas, and the production of stars of intermediate mass stops within a million years or so of the formation of the highest mass stars. Finally, we consider what we know directly and indirectly about the IMF of the more distant members of the Local Group: NGC 6822, IC 1613, M 31, M 33, and IC 10. Of these, only the “starburst” galaxy IC 10 shows signs of anything odd in its massive star population.

To appear in: The Stellar Initial Mass Function, proceedings of the 38th Herstmonceux Conference, ed. G. Gilmore, I. Parry, and S. Ryan For preprints, contact massey@noao.edu

Also available from the URL ftp://tofu.tuc.noao.edu/pub/cam.ps
or by anonymous ftp at tofu.tuc.noao.edu; cd pub; get cam.ps

Meeting Announcements

IAU Symposium 192: The Stellar Content of Local Group Galaxies

Patricia Whitelock^{1} and Russell Cannon^{2}

1 SAAO, Cape Town, South Africa; Chair of LOC; paw@da.saao.ac.za 2 AAO, Sydney, Australia; Chair of SOC; rdc@aaoepp.aao.gov.au

IAU Symposium 192 will be held in Cape Town, South Africa, from 7–11 September 1998. The rather specific topic of the meeting is to be “The Stellar Content of Local Group Galaxies”. The objective is to pull together the latest data on stars in Local Group galaxies. These include accurate colour-magnitude diagrams from HST and from wide-field CCD imagers on medium-sized groundbased telescopes, and spectroscopic data for faint stars from multi-object systems or high-dispersion spectrographs on the largest telescopes. The emphasis is to be on understanding the stellar content of individual galaxies in the Local Group, with its implications for the formation and evolution of galaxies. A second major objective will be the application of the results to more distant systems, where individual stars can no longer be observed.

We are liaising with the organisers of the IAU Symposium 190 on the Magellanic Clouds, to be held in Victoria BC in July 1998, to ensure that the two scientific programmes are complementary.

For further information, registration forms etc., see the WWW page at the following URL:

http://www.saao.ac.za/~lgroup, or send e-mail to lgroup@saao.ac.za, or if really stuck, contact

one of us directly.